The safe stopping distance, d, in metres, for a boat travelling at v km/h in calm water is determined to be d = 0.002 (2v^2 + 10v + 3000) .
a) What is the safe stopping distance if the speed is 12 km/h?
b) What is the initial speed of the boat if it takes 15 m to stop?

Answer by Boreal(15235)

if speed is 12 km/h
=0.002[2*12^2+10(12)+3000)]
=6.528 m
-
15=.002(2v^2+10v+3000)=0.004v^2+0.02v+6
so 0.004v^2+0.02c-9=0
4v^2+20v-9000=0
v^2+5v-2250=0
(v+50)(v-45)=0
v=45 km/h